# Analysis of Clustering and Degree Index in Random Graphs and Complex Networks

### v1

@@ 3.2 Further analysis of $\mathbb{E}[\mathrm{CI}_{2}(\mathcal{G})]$
Now that we know $\mathbb{E}[\mathrm{CI}_{2}(\mathcal{G})]$ is sublinear via Corollary 3.1 when $\mathcal{G}$ is an Erdős-Rényi graph, we will focus on this expectation in more detail and provide a more precise analysis for this case. This will require an understanding of expectations of the form $\mathbb{E}[C(i)C(j)]$ for nodes $i\neq j$.
@@
for each $n$, where $D_{2}$ is a positive constant depending on $p$, but not on $n$
@@
$\leq n(n-1)\left(\mathbb{E}[C(i)^{2}]-\mathbb{E}[C(i)C(j)]\right)\leq n^{2}\left(p^{2}+\frac{D_{1}}{n^{2}}-p^{2}+\frac{D_{2}}{n^{2}}\right)\leq D_{1}+D_{2}=:K.$
@@
Proof of Proposition 3.2. Let $A_{i}$ be the event that $|d_{i}-(n-1)p|<n^{2/3}$. Define $A_{j}$ similarly. Then,
@@
where $S$ consists of small terms that are easy to estimate. Indeed, as in the proof of Proposition 3.1, via McDiarmid inequality, we may reach at the conclusion that we can bound $S$ by $\frac{D}{n^{2}}$ for some constant $D$ independent of $n$. (Again $D$’s in distinct appearances may denote different constants.) So let us focus on $\mathbb{E}[C(i)C(j)\mathbf{1}(A_{i})\mathbf{1}(A_{j})]$.
Let $M$ be the event that the edge between $i,j$ exists. Also, let $T_{i}$ and $T_{j}$ denote the number of triangles containing $i$ and $j$ as a node, respectively. Then
@@
Now we will examine the conditional expectation of $T_{i}T_{j}$ given $d_{i},d_{j},M$ and again of $T_{i}T_{j}$ given $d_{i},d_{j},\overline{M}$ in more detail. First, we decompose $T_{i}$ and $T_{j}$ as $T_{i1}+T_{i2}+T_{i3}$ and $T_{j1}+T_{j2}+T_{j3}$ as follows. $T_{i1}=T_{j1}$ denotes the number of triangles that have both $i$ and $j$ as vertices. $T_{i2}=T_{j2}$ denotes the number of edges $e_{k\ell}$, with $k$ and $\ell$ distinct from $i$ and $j$, and the triangles $\{i,k,\ell\}$ and $\{j,k,\ell\}$ both appear in the graph. $T_{i3}$ and $T_{j3}$ denote the remaining triangles containing $i$ and $j$ as a node, respectively.
@@
$\mathbb{E}[T_{i}T_{j}\mid\overline{M},d_{i},d_{j}]=\mathbb{E}[T_{i}\mid\overline{M},d_{i},d_{j}]\mathbb{E}[T_{j}\mid\overline{M},d_{i},d_{j}]+\mathbb{E}[T_{i2}^{2}-(\mathbb{E}[T_{i2}|\overline{M},d_{i},d_{j}])^{2}|\overline{M},d_{i},d_{j}].$
Substituting these into our previous formula for $\mathbb{E}[C(i)C(j)\mathbf{1}(A_{i})\mathbf{1}(A_{j})]$, noting that
$\mathbb{E}[T_{i}\mid M,d_{i},d_{j}]=\mathbb{E}[T_{i}\mid\overline{M},d_{i},d_{j}]=p^{2}\binom{d_{i}}{2},$
and that a similar result holds if we replace $T_{i}$ by $T_{j}$, we obtain,
@@
Also, since $\mathbb{P}(\overline{A_{i}})\leq M_{1}e^{-M_{2}n^{1/3}}$ for some constants $M_{1},M_{2}$ independent of $n$, and since a similar bound holds for $A_{j}$, we have $\mathbb{P}(A_{i}\cap A_{j})\geq 1-\frac{D}{n^{2}}$ for some constant $D$ independent of $n$. Using this, we reach at,
@@ 4 Degree index in Erdős-Rényi graphs
As already noted earlier, the degree index (or, total degree irregularity) is previously introduced and studied, but we are not aware of a relevant analysis for random graphs. This section is devoted to such an analysis, and this turns out to be more tractable compared to the clustering index case due to the lack of $\binom{d_{i}}{2}$ terms in the denominator.
We begin the discussion with the $\alpha=2$ case. Let $\mathcal{G}$ be an Erdős-Rényi graph with parameters $n\in\mathbb{N}$ and $p\in(0,1)$. We are interested in calculating
@@
Let $k=n-2$ for convenience. Since $d_{i}\sim\mathrm{Bin}(n-1,p)=\mathrm{Bin}(k+1,p)$, we have
$\mathbb{E}[d_{i}^{2}]=\mathbb{E}[d_{i}]^{2}+\mathrm{Var}(d_{i})=(n-1)^{2}p^{2}+(n-1)p(1-p)=(k+1)^{2}p^{2}+(k+1)p(1-p)$
Next, let us examine $\mathbb{E}[d_{i}d_{j}]$. Let $A$ denote the event that $e_{ij}$ appears in the graph. Conditioning on $A$ and $\overline{A}$ yields:
$\mathbb{E}[d_{i}d_{j}]=p\mathbb{E}[d_{i}d_{j}\mid A]+(1-p)\mathbb{E}[d_{i}d_{j}\mid\overline{A}]$
If we assume $e_{ij}$ does not appear in the graph, then the number of possible neighbors of $i$ or $j$ is now $n-2$ instead of $n-1$. And since we are in the Erdős-Renyi model, the presence of any other edge is independent of the presence of $e_{ij}$. Moreover, apart from whether $e_{ij}$ appears or not, $d_{i}$ and $d_{j}$ are independent. By the above observations, $d_{i}\mid A$ and $d_{j}\mid A$ are i.i.d. with the common distribution $1+\mathrm{Bin}(n-2,p)=1+\mathrm{Bin}(k,p)$. Similarly, $d_{i}\mid\overline{A}$ and $d_{j}\mid\overline{A}$ are i.i.d. this time with the common distribution $\mathrm{Bin}(n-2,p)=\mathrm{Bin}(k,p)$. Therefore, we have
$\mathbb{E}[d_{i}d_{j}]$ $=p\mathbb{E}[d_{i}d_{j}\mid A]+(1-p)\mathbb{E}[d_{i}d_{j}\mid\overline{A}]=p(kp+1)^{2}+(1-p)(kp)^{2}$
@@
Recalling that $\mathrm{DI}_{2}(\mathcal{G})=\sum_{1\leq i<j\leq n}|d_{i}-d_{j}|^{2}$, we conclude the following.
Theorem 4.1
$\mathbb{E}[\mathrm{DI}_{2}(\mathcal{G})]=6\binom{n}{3}p(1-p).$
In particular,
$\frac{\mathbb{E}[\mathrm{DI}_{2}(\mathcal{G})]}{6\binom{n}{3}}=p(1-p),$
which is independent of $n$.
Next we focus on $\mathrm{DI}_{1}(\mathcal{G})$. In this case a direct computation readily gives
@@
$\mathbb{E}[|d_{i}-d_{j}|]=p\mathbb{E}[|d_{i}-d_{j}|\mid A]+(1-p)\mathbb{E}[|d_{i}-d_{j}|\mid\overline{A}].$
Note that the distribution of $d_{i}-d_{j}$ is the same regardless of whether $e_{ij}$ is present in the graph or not. And since the only dependence between $d_{i}$ and $d_{j}$ was a result of $e_{ij}$, we have,
@@
where $B_{1},B_{2}$ are independent binomial random variables with parameters $n-2$ and $p$. For general $p$ case, an easy upper bound on $\mathbb{E}[|d_{i}-d_{j}|]$ can be obtained as follows:
@@
Improving this, we will show an exact asymptotic result for $\mathbb{E}[\mathrm{DI}_{1}(\mathcal{G})]$. This requires the following result on the mean absolute difference of binomials.
Proposition 4.1
@@
This proposition, whose proof is given at the end of this section, immediately yields the following result on $\mathrm{DI}_{1}$.
@@
$\mathbb{E}[\mathrm{DI}_{1}(\mathcal{G})]=\frac{(n-2)\binom{2n-4}{n-2}}{2^{2n-4}}.$
In particular, $\mathbb{E}[\mathrm{DI}_{1}(\mathcal{G})]\sim\sqrt{\frac{n-2}{\pi}}$, as $n\rightarrow\infty$.
@@
$\mathbb{E}[\mathrm{DI}_{1}(\mathcal{G})]\sim\frac{2}{\sqrt{\pi}}\sqrt{(n-2)p(1-p)}$

### v2

@@ 3.2 Further analysis of $\mathbb{E}[\mathrm{CI}_{2}(\mathcal{G})]$
Now that we know $\mathbb{E}[\mathrm{CI}_{2}(\mathcal{G})]$ is sublinear via Corollary 3.1 when $\mathcal{G}$ is an Erdős-Rényi graph, we will focus on this expectation in more detail and provide a more precise analysis for this case. This will require understanding of expectations of the form $\mathbb{E}[C(i)C(j)]$ for nodes $i\neq j$.
@@
for each $n$, where $D_{2}$ is a positive constant depending on $p$, but not on $n$.
@@
$=n(n-1)\left(\mathbb{E}[C(i)^{2}]-\mathbb{E}[C(i)C(j)]\right)\leq n^{2}\left(p^{2}+\frac{D_{1}}{n^{2}}-p^{2}+\frac{D_{2}}{n^{2}}\right)=D_{1}+D_{2}=:K.$
@@
Proof of Proposition 3.2. Recall from the proof of Proposition 3.1 that $A_{i}$ is the event that $|d_{i}-(n-1)p|<n^{2/3}$. Define $A_{j}$ similarly. We have
@@
where
$S=\mathbb{E}[C(i)C(j)\mathbf{1}(A_{i})\mathbf{1}(\overline{A_{j}})]+\mathbb{E}[C(i)C(j)\mathbf{1}(\overline{A_{i}})\mathbf{1}(A_{j})]+\mathbb{E}[C(i)C(j)\mathbf{1}(\overline{A_{i}})\mathbf{1}(\overline{A_{j}})].$
Noting that $A_{i}$ and $A_{j}$ are equally likely, and using trivial upper bound for the clustering coefficients, we see that
$S\leq 3\mathbb{P}(\overline{A_{i}}).$
Recalling now $\mathbb{P}(\overline{A_{i}})\leq M_{1}e^{-M_{2}n^{1/3}}$ for some constants $M_{1},M_{2}>0$ from the Proof of Proposition 3.1, we conclude that
$S\leq\frac{D}{n^{2}},\quad n\geq 1,$
where $D$ is a positive constant independent of $n$. (Again, $D$’s in distinct appearances may denote different constants.) Next, we focus on the estimation of $\mathbb{E}[C(i)C(j)\mathbf{1}(A_{i})\mathbf{1}(A_{j})]$.
Let $M$ be the event that the edge between $i,j$ exists. Also, let $T_{i}$ and $T_{j}$ denote the number of triangles that contain $i$ and $j$ as a node, respectively. Then
@@
Now we will examine the conditional expectation of $T_{i}T_{j}$ given $d_{i},d_{j},M$ and again of $T_{i}T_{j}$ given $d_{i},d_{j},\overline{M}$ in more detail. For this purpose, we decompose $T_{i}$ and $T_{j}$ as
$T_{i}=T_{i1}+T_{i2}+T_{i3}\quad\text{and}\quad T_{j}=T_{j1}+T_{j2}+T_{j3},$
where,
  • $T_{i1}=T_{j1}$ denotes the number of triangles that have both $i$ and $j$ as vertices,
  • $T_{i2}=T_{j2}$ denotes the number of edges $e_{k\ell}$, with $k$ and $\ell$ distinct from $i$ and $j$, and the triangles $\{i,k,\ell\}$ and $\{j,k,\ell\}$ both appear in the graph,
  • $T_{i3}$ and $T_{j3}$ denote the remaining triangles containing $i$ and $j$ as a node, respectively.
@@
$\mathbb{E}[T_{i}T_{j}\mid\overline{M},d_{i},d_{j}]=\mathbb{E}[T_{i}\mid\overline{M},d_{i},d_{j}]\mathbb{E}[T_{j}\mid\overline{M},d_{i},d_{j}]+\mathbb{E}[T_{i2}^{2}-(\mathbb{E}[T_{i2}|\overline{M},d_{i},d_{j}])^{2}|\overline{M},d_{i},d_{j}].$ (7)
Now, note that
$\mathbb{E}[T_{i}\mid M,d_{i},d_{j}]=\mathbb{E}[T_{i}\mid\overline{M},d_{i},d_{j}]=p\binom{d_{i}}{2}$. (8)
To see this, given the degree $d_{i}$ of $i$th node along with $d_{j}$ and $M$, observe that there are $\binom{d_{i}}{2}$ possible triangles containing node $i$, and one just needs to count the number of edges between the pairs of neighbors of $i$ to find $T_{i}$. Hence, there are $\binom{d_{i}}{2}$ candidates for triangles containing node $i$, each of which is present with probability $p$, independent of others. Thus, one has
$\mathbb{E}[T_{i}\mid M,d_{i},d_{j}]=p\binom{d_{i}}{2}$.
Noting that given the degree of $d_{i}$, the presence or absence of the edge between node $i$ and $j$ does not change the distribution of $T_{i}$, we similarly have
$\mathbb{E}[T_{i}\mid\overline{M},d_{i},d_{j}]=p\binom{d_{i}}{2}$.
Combining our observations in (3.2), (3.2), (7) and (8), we obtain,
@@
Also, again recalling from the proof of Proposition 3.1 that $\mathbb{P}(\overline{A_{i}})\leq M_{1}e^{-M_{2}n^{1/3}}$ for some constants $M_{1},M_{2}$ independent of $n$, and noting a similar bound holds for $A_{j}$, we have $\mathbb{P}(A_{i}\cap A_{j})\geq 1-\frac{D}{n^{2}}$ for some constant $D$ independent of $n$. Using this, we reach at,
@@ 4 Degree index in Erdős-Rényi graphs
As already noted earlier, the degree index (or total degree irregularity) has been previously introduced and studied, but we are not aware of a relevant analysis for random graphs. This section is devoted to such an analysis, and this turns out to be more tractable compared to the clustering index case due to the lack of $\binom{d_{i}}{2}$ terms in the denominator. We begin the discussion by studying $\mathbb{E}[\mathrm{DI}_{2}(\mathcal{G})]$.
Theorem 4.1
$\mathbb{E}[\mathrm{DI}_{2}(\mathcal{G})]=6\binom{n}{3}p(1-p).$
In particular,
$\frac{\mathbb{E}[\mathrm{DI}_{2}(\mathcal{G})]}{6\binom{n}{3}}=p(1-p),$
which is independent of $n$.
Proof. Let $\mathcal{G}$ be an Erdős-Rényi graph with parameters $n\in\mathbb{N}$ and $p\in(0,1)$. We are interested in calculating
@@
Let k=n2k=n-2 for convenience. Denoting the binomial distribution by $\mathrm{Bin}$, since $d_{i}\sim\mathrm{Bin}(n-1,p)=\mathrm{Bin}(k+1,p)$, we have
$\mathbb{E}[d_{i}^{2}]=(\mathbb{E}[d_{i}])^{2}+\mathrm{Var}(d_{i})=(n-1)^{2}p^{2}+(n-1)p(1-p)=(k+1)^{2}p^{2}+(k+1)p(1-p).$
Next, let us examine $\mathbb{E}[d_{i}d_{j}]$. Let $M$ denote the event that $e_{ij}$ appears in the graph. Conditioning on $M$ and $\overline{M}$ yields
$\mathbb{E}[d_{i}d_{j}]=p\mathbb{E}[d_{i}d_{j}\mid M]+(1-p)\mathbb{E}[d_{i}d_{j}\mid\overline{M}].$
If we assume that $e_{ij}$ does not appear in the graph, then the number of possible neighbors of $i$ or $j$ is now n2n-2 instead of n1n-1. And since we are in the Erdős-Rényi model, the presence of any other edge is independent of the presence of $e_{ij}$. Moreover, apart from whether $e_{ij}$ appears or not, $d_{i}$ and $d_{j}$ are independent. By the above observations, $d_{i}\mid M$ and $d_{j}\mid M$ are i.i.d. with the common distribution $1+\mathrm{Bin}(n-2,p)=1+\mathrm{Bin}(k,p)$. Similarly, $d_{i}\mid\overline{M}$ and $d_{j}\mid\overline{M}$ are i.i.d. this time with the common distribution $\mathrm{Bin}(n-2,p)=\mathrm{Bin}(k,p)$. Therefore, we have
$\mathbb{E}[d_{i}d_{j}]$ $=p\mathbb{E}[d_{i}d_{j}\mid M]+(1-p)\mathbb{E}[d_{i}d_{j}\mid\overline{M}]=p(kp+1)^{2}+(1-p)(kp)^{2}$
@@
Recalling that $\mathrm{DI}_{2}(\mathcal{G})=\sum_{1\leq i<j\leq n}|d_{i}-d_{j}|^{2}$, result follows. $\square$
Next, we focus on $\mathrm{DI}_{1}(\mathcal{G})$. In this case a direct computation readily gives
@@
$\mathbb{E}[|d_{i}-d_{j}|]=p\mathbb{E}[|d_{i}-d_{j}|\mid M]+(1-p)\mathbb{E}[|d_{i}-d_{j}|\mid\overline{M}].$
Note that the distribution of $d_{i}-d_{j}$ is the same regardless of whether the edge between $i$ and $j$, $e_{ij}$, is present in the graph or not. And since the only dependence between $d_{i}$ and $d_{j}$ was a result of $e_{ij}$, we have,
@@
where $B_{1},B_{2}$ are independent binomial random variables with parameters n2n-2 and $p$. For general $p$ case, an easy upper bound on $\mathbb{E}[|d_{i}-d_{j}|]$ can be obtained as follows:
@@
Improving this, one may obtain an exact asymptotic result for $\mathbb{E}[\mathrm{DI}_{1}(\mathcal{G})]$. This will follow from the following elementary result on the mean absolute difference of binomials.
Lemma 4.1
@@
Although the lemma is elementary and there are similar results in the literature (see, for example, [9]), we were unable to find the exact form used in our case. Therefore, we include the proof of the lemma in the appendix.
Now, given Lemma 4.1, the following result on $\mathrm{DI}_{1}$ follows immediately.
@@
$\mathbb{E}[\mathrm{DI}_{1}(\mathcal{G})]=\frac{(n-2)\binom{n}{2}\binom{2n-4}{n-2}}{2^{2n-4}}.$
In particular, $\mathbb{E}[\mathrm{DI}_{1}(\mathcal{G})]\sim\sqrt{\frac{n-2}{\pi}}\binom{n}{2}$, as $n\rightarrow\infty$.
@@
$\mathbb{E}[\mathrm{DI}_{1}(\mathcal{G})]\sim\frac{2}{\sqrt{\pi}}\binom{n}{2}\sqrt{(n-2)p(1-p)}$